# Functors on the category of finite sets revisited

### v2

@@ 5. Projectives in characteristic zero
The purpose of this section is to introduce the indecomposable projectives of the category $\mathcal{F}(\mathbf{FA})$. Apart from the projective $P^{\mathbf{FA}}_{\mathbf{0}}\cong\mathbbm{k}$, these are indexed by partitions $\lambda\vdash n$ for $0<n$. The corresponding classification of the simple functors in $\mathcal{F}(\mathbf{FA})$ is treated in Section 6. The case of the partitions $(1^{n})$ is exceptional, accounting for the subtleties of the theory.
Theorem 5.4 introduces the objects $S_{\lambda}(\overline{P})$; these are projective for $\lambda$ not of the form $(1^{n})$. For $\lambda=(1^{n})$, $S_{\lambda}(\overline{P})=\Lambda^{n}(\overline{P})$ is not projective; its projective cover is identified as $\Lambda^{n+1}(P^{\mathbf{FA}})$ in Theorem 5.10, which also describes the structure of this projective. (The latter results are contained in the work of Wiltshire-Gordon [Wil14].)
@@
The general result Corollary 4.7 implies that the indecomposable projective objects in $\mathcal{F}(\mathbf{FA})$ have finite length.
Notation 5.2.
Write $P^{\mathbf{FA}}$ as shorthand for $P^{\mathbf{FA}}_{\mathbf{1}}$.
@@ 5.1. A family of indecomposable projectives
$\mathbbm{k}\mathfrak{S}_{t}\cong\bigoplus_{\lambda\vdash t}S_{\lambda}\boxtimes S_{\lambda},$
@@
For $t\in\mathbb{N}$, the decomposition $(\overline{P})^{\otimes t}\cong\bigoplus_{\lambda\vdash t}S_{\lambda}(\overline{P})^{\oplus\dim S_{\lambda}}$ is the complete splitting of $(\overline{P})^{\otimes t}$ into indecomposable objects. Moreover, for partitions $\lambda,\mu\vdash t$,
@@
By Theorem 4.4, the endomorphism ring of $(\overline{P})^{\otimes t}$ in $\mathcal{F}(\mathbf{FA})$ is $\mathbbm{k}\mathfrak{S}_{t}$. The result follows by standard arguments. ∎
@@
To prove that $S^{2}(\overline{P})$ is projective, it suffices to define a suitable idempotent in $\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{2}})$ inducing a projection onto $S^{2}(\overline{P})$. Now, $\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{2}})$ identifies as a $\mathbbm{k}$-vector space as $P^{\mathbf{FA}}_{\mathbf{2}}(\mathbf{2})$, by Yoneda. Using this isomorphism, one checks that 12(([$-\frac{1}{2}\big{(}([1]-[2])\otimes([1]-[2])\big{)}\in P^{\mathbf{FA}}_{\mathbf{2}}(\mathbf{2})$ is such an idempotent.
@@
    An alternative proof of the Theorem is based on Lemma 5.9 below. In particular, this exhibits a surjection $P^{\mathbf{FA}}\otimes\overline{P}\twoheadrightarrow\Lambda^{2}(P^{\mathbf{FA}})$ between projective functors, using Corollary 3.10 for the projectivity of the domain. The kernel identifies with $S^{2}(\overline{P})$, which is thus projective, since it is a direct summand of the projective $P^{\mathbf{FA}}\otimes\overline{P}$, by the splitting of the short exact sequence. By induction, this argument generalizes using Lemma 5.9 to prove the Theorem (or one can proceed as above).
@@
    This result was contained in a preliminary version of [PV23], where it was proved by exploiting the homological consequences of the results of that paper.
@@ 5.2. The structure of $\Lambda^{t}(P^{\mathbf{FA}})$
In this section we consider the projective $\Lambda^{t}(P^{\mathbf{FA}})$ and its subobject $\Lambda^{t}(\overline{P})$ for $0<t\in\mathbb{N}$. This is related to the analysis given by Wiltshire-Gordon [Wil14] (who does not use $\overline{P}$).
Convention 5.6.
@@
By definition, $\Lambda^{t}(P^{\mathbf{FA}})(\mathbf{t})\cong\Lambda^{t}(\mathbbm{k}\mathbf{t})$ as a $\mathbbm{k}\mathfrak{S}_{t}$-module, and the latter identifies with $\mathsf{sgn}_{t}$. For the case $\Lambda^{t}(\overline{P})(\mathbf{t+1})$, it suffices to show that, if $t>0$, the action of $\mathfrak{S}_{t+1}$ is non-trivial. This is checked by considering the action of the transposition $(12)$ and using the basis of $\overline{P}(\mathbf{t+1})$ given by the non-zero elements of the form $[i]-[t+1]$. ∎
@@
For $t\in\mathbb{N}$, by convention, $\Lambda^{0}(\overline{P})=\overline{\mathbbm{k}}$, which is simple. Henceforth we suppose that $t>0$. By construction, $\Lambda^{t}(\overline{P})$ is a direct summand of $\overline{P}^{\otimes t}$. Hence, any composition factor of $\Lambda^{t}(\overline{P})$ is also one of $\overline{P}^{\otimes t}$. Now, Proposition 4.6 implies that any such composition factor is non-zero when evaluated on $\mathbf{t+1}$. Since $\dim\Lambda^{t}(\overline{P})(\mathbf{t+1})=1$ by Lemma 5.7, it follows that there is only one such. ∎
@@
$[x_{1}]\otimes\ldots\otimes[x_{t}]\mapsto[x_{1}]\wedge\ldots\wedge[x_{t}].$
@@
For the surjectivity, it suffices to observe that the image of $[x_{1}]\otimes([x_{2}]-[x_{1}])\otimes\ldots\otimes([x_{t}]-[x_{1}])$ is $[x_{1}]\wedge\ldots\wedge[x_{t}]$. The final statement holds since $\Lambda^{t}(P^{\mathbf{FA}})$ is projective. ∎
@@
To show that this exhibits $\Lambda^{t}(P^{\mathbf{FA}})$ as the projective cover of the simple functor $\Lambda^{t-1}(\overline{P})$, it suffices to show that $\Lambda^{t}(P^{\mathbf{FA}})$ is indecomposable. For $t=0$ this is Lemma 3.3, hence we suppose that $t>0$. It suffices to show that $\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{t}(P^{\mathbf{FA}}))=\mathbbm{k}$. Using the canonical surjection $P^{\mathbf{FA}}_{\mathbf{t}}\twoheadrightarrow\Lambda^{t}(P^{\mathbf{FA}})$, this follows from the equality $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{t}},\Lambda^{t}(P^{\mathbf{FA}}))=\mathbbm{k}$ given by Yoneda’s lemma and the fact that that $\Lambda^{t}(\mathbbm{k}[\mathbf{t}])\cong\mathbbm{k}$, by Lemma 5.7. ∎
The short exact sequences (5.1) splice to yield the following (a related result occurs in [Wil14]):
@@
$\ldots\rightarrow\Lambda^{t+1}(P^{\mathbf{FA}})\rightarrow\Lambda^{t}(P^{\mathbf{FA}})\rightarrow\Lambda^{t-1}(P^{\mathbf{FA}})\rightarrow\ldots\rightarrow\Lambda^{1}(P^{\mathbf{FA}})\rightarrow\Lambda^{0}(P^{\mathbf{FA}})=\mathbbm{k}\rightarrow\mathbbm{k}_{\mathbf{0}}\rightarrow 0.$
@@
For $t$ a positive integer, Wiltshire-Gordon [Wil14] defines $D_{t}$ to be the cokernel of the map $\Lambda^{t+1}(P^{\mathbf{FA}})\rightarrow\Lambda^{t}(P^{\mathbf{FA}})$ appearing in Proposition 5.11. By the above identifications, one has $D_{t}\cong\Lambda^{t-1}(\overline{P})$.
@@ 5.3. Classification of the indecomposable projectives in $\mathcal{F}(\mathbf{FA})$
For $n\in\mathbb{N}$, Theorem 5.4 gives that $S_{\lambda}(\overline{P})$ is an indecomposable projective for $\lambda\neq(1^{n})$ and Theorem 9.3 that $\Lambda^{n}(P^{\mathbf{FA}})$ is an indecomposable projective. One checks using Proposition 5.3 that these are pairwise non-isomorphic, in the sense of the unicity in the statement. It remains to show that these exhaust the isomorphism classes of indecomposable projectives.
@@
The maximality of $n$ implies that either $Q$ is isomorphic to $\Lambda^{n+1}(P^{\mathbf{FA}})$ or $Q$ is a quotient of $(\overline{P}^{\otimes n+1}/\Lambda^{n+1}(\overline{P}))$ (these possibilities are mutually exclusive). In the second case, using the splitting of $\overline{P}^{\otimes n+1}/\Lambda^{n+1}(\overline{P})$ into isotypical components, $Q$ must be isomorphic to $S_{\mu}(\overline{P})$, for a unique $\mu\vdash n+1$ such that $\mu\neq(1^{n+1})$. ∎
@@
This theorem provides an abstract classification of the simple functors in $\mathcal{F}(\mathbf{FA})$. Such a classification was given by Wiltshire-Gordon in [Wil14], using a slightly different method. The simple functors are made explicit in Section 6 below.
@@ 6. The simple $\mathbbm{k}\mathbf{FA}$-modules
In this section, the simple objects of $\mathcal{F}(\mathbf{FA})$ (aka. $\mathbbm{k}\mathbf{FA}$-modules) are made explicit. These are related to the structure of $\mathbbm{k}{\mathbf{FI}}(-,-)$, thus relating to results of Feigin and Shoikhet [FS07].
@@ 6.1. The simple functors
Consider the case $n>0$; Corollary 3.11 exhibits $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$ as a quotient of $P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n-1}$. As in the proof of Theorem 5.14, the latter decomposes as $\Lambda^{n}(P^{\mathbf{FA}})\oplus(\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P}))\oplus(\overline{P}^{\otimes n-1}/\Lambda^{n-1}(\overline{P}))$, in which each summand is projective. Moreover, $\overline{P}^{\otimes n-1}/\Lambda^{n-1}(\overline{P})$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n-1}}$. Since $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$ is zero when evaluated on $\mathbf{n-1}$, Yoneda’s lemma implies that there are no non-trivial maps from $\overline{P}^{\otimes n-1}/\Lambda^{n-1}(\overline{P})$ to $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$. Hence the surjection of Corollary 3.11 restricts to a surjection
@@
To conclude, it suffices to check that this restriction is not surjective. This is seen by evaluating on $\mathbf{n}$. ∎
@@
We first prove that, for $\lambda\neq(1^{n})$, $C_{\lambda}=\tilde{C}_{\lambda}$ is simple. Suppose otherwise, since $C_{\lambda}(\mathbf{n})$ is a simple $\mathfrak{S}_{n}$-module by Lemma 6.5, then $C_{\lambda}$ would have a non-zero composition factor $G$ such that $G(\mathbf{n})=0$. Since $C_{\lambda}$ is a quotient of $\overline{P}^{\otimes n}$ by Corollary 6.3, $G$ is a composition factor of the latter, hence Proposition 4.6 gives that $G(\mathbf{n+1})\neq 0$.
Now, our understanding of the indecomposable projective generators of $\mathcal{F}(\mathbf{FA})$ (see Theorem 5.14) implies that either $G$ is a quotient of $S_{\mu}(\overline{P})$ for some partition $\mu\vdash n+1$ or $G$ is $\Lambda^{n}(\overline{P})$, hence a quotient of $\Lambda^{n+1}(P^{\mathbf{FA}})$. The first case is excluded since $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(S_{\mu}(\overline{P}),\overline{P}^{\otimes n})=0$, as a consequence of Theorem 4.4, since $S_{\mu}(\overline{P})$ is a direct summand of $\overline{P}^{\otimes n+1}$. The second case is excluded similarly, since $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n+1}(P^{\mathbf{FA}}),S_{\lambda}(\overline{P}))=0$ for $\lambda\neq(1^{n})$, again as a consequence of Theorem 4.4; explicitly, any such morphism would factor across $\Lambda^{n}(\overline{P})$ and $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n}(\overline{P}),S_{\lambda}(\overline{P}))=0$ by Proposition 5.3.
@@
In [SST22], Sam, Snowden and Tosteson gave a model for the category of $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules using polynomial representations of the Witt Lie algebra. They observed that, under their equivalence (and using duality), the classification of the simple $\mathbbm{k}\mathbf{FA}$-modules given in [Wil14] corresponds to the classification of certain irreducible modules over the Witt Lie algebras given by Feigin and Shoikhet [FS07].

### v3

@@ 5. Projectives in characteristic zero
In this section, we study the indecomposable projectives of the category $\mathcal{F}(\mathbf{FA})$. Apart from the projective $P^{\mathbf{FA}}_{\mathbf{0}}\cong\mathbbm{k}$, these are indexed by partitions $\lambda\vdash n$ for $0<n$. The corresponding classification of the simple functors in $\mathcal{F}(\mathbf{FA})$ is treated in Section 6. The case of the partitions $(1^{n})$ is exceptional, accounting for the subtleties of the theory.
Theorem 5.3 introduces the objects $S_{\lambda}(\overline{P})$; these are projective for $\lambda$ not of the form $(1^{n})$. For $\lambda=(1^{n})$, $S_{\lambda}(\overline{P})=\Lambda^{n}(\overline{P})$ is not projective, where $P^{\mathbf{FA}}$ is shorthand for $P^{\mathbf{FA}}_{\mathbf{1}}$; its projective cover is identified as $\Lambda^{n+1}(P^{\mathbf{FA}})$ in Theorem 5.9, which also describes the structure of this projective. (The latter results are contained in the work of Wiltshire-Gordon [10].)
@@
Recall that Corollary 4.33 established that the indecomposable projective objects in $\mathcal{F}(\mathbf{FA})$ have (finite length) composition series.
@@ 5.1. A family of indecomposable projectives
(5.1) $\mathbbm{k}\mathfrak{S}_{t}\cong\bigoplus_{\lambda\vdash t}S_{\lambda}\boxtimes S_{\lambda},$
@@
For $t\in\mathbb{N}$, the decomposition $\overline{P}^{\otimes t}\cong\bigoplus_{\lambda\vdash t}S_{\lambda}(\overline{P})^{\oplus\dim S_{\lambda}}$ is the complete splitting of $\overline{P}^{\otimes t}$ into indecomposable objects. Moreover, for partitions $\lambda,\mu\vdash t$,
@@
By Corollary 4.21, the endomorphism ring of $\overline{P}^{\otimes t}$ in $\mathcal{F}(\mathbf{FA})$ is $\mathbbm{k}\mathfrak{S}_{t}$. Let $e_{\lambda}$ and $e_{\mu}$ be idempotents in $\mathbbm{k}\mathfrak{S}_{t}$ so that there are isomorphisms of $\mathbbm{k}\mathfrak{S}_{t}$-modules $S_{\lambda}\cong\mathbbm{k}\mathfrak{S}_{t}e_{\lambda}$ and $S_{\mu}\cong\mathbbm{k}\mathfrak{S}_{t}e_{\mu}$. Thus there are isomorphisms $S_{\lambda}(\overline{P})\cong\overline{P}^{\otimes t}e_{\lambda}$ and $S_{\mu}(\overline{P})\cong\overline{P}^{\otimes t}e_{\mu}$ so that one has the isomorphism of $\mathbbm{k}$-vector spaces
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(S_{\lambda}(\overline{P}),S_{\mu}(\overline{P}))\cong e_{\lambda}\mathbbm{k}\mathfrak{S}_{t}e_{\mu}.$
Using the identification of $\mathbbm{k}\mathfrak{S}_{t}$ as a bimodule given by (5.1), the result follows. ∎
@@
To prove that $S^{2}(\overline{P})$ is projective, it suffices to define a suitable idempotent in $\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{2}})$ inducing a projection onto $S^{2}(\overline{P})$. Now, $\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{2}})$ identifies as a $\mathbbm{k}$-vector space as $P^{\mathbf{FA}}_{\mathbf{2}}(\mathbf{2})$, by Yoneda. Using this isomorphism, one checks that $\frac{1}{2}\big{(}([1]-[2])\otimes([1]-[2])\big{)}\in P^{\mathbf{FA}}_{\mathbf{2}}(\mathbf{2})$ is such an idempotent.
@@
    An alternative proof of the Theorem is based on Lemma 5.8 below. In particular, this exhibits a surjection $P^{\mathbf{FA}}\otimes\overline{P}\twoheadrightarrow\Lambda^{2}(P^{\mathbf{FA}})$ between projective functors, using Proposition 3.10 for the projectivity of the domain. The kernel identifies with $S^{2}(\overline{P})$, which is thus projective, since it is a direct summand of the projective $P^{\mathbf{FA}}\otimes\overline{P}$, by the splitting of the short exact sequence. By induction, this argument generalizes using Lemma 5.8 to prove the Theorem (or one can proceed as above).
@@
    This result was contained in a preliminary version of [7], where it was proved by exploiting the homological consequences of the results of that paper.
@@ 5.2. The structure of $\Lambda^{t}(P^{\mathbf{FA}})$
In this section we consider the projective $\Lambda^{t}(P^{\mathbf{FA}})$ and its subobject $\Lambda^{t}(\overline{P})$ for $0<t\in\mathbb{N}$. This is related to the analysis given by Wiltshire-Gordon [10] (who does not use $\overline{P}$).
Convention 5.5.
@@
By definition, $\Lambda^{t}(P^{\mathbf{FA}})(\mathbf{t})\cong\Lambda^{t}(\mathbbm{k}\mathbf{t})$ as a $\mathbbm{k}\mathfrak{S}_{t}$-module, and the latter identifies with $\mathsf{sgn}_{t}$. For the case $\Lambda^{t}(\overline{P})(\mathbf{t+1})$, it suffices to show that, if $t>0$, the action of $\mathfrak{S}_{t+1}$ is non-trivial. This is checked by considering the action of the transposition $(12)$ and using the basis of $\overline{P}(\mathbf{t+1})$ given by the non-zero elements of the form [i][t+1][i]-[t+1]. ∎
@@
For $t\in\mathbb{N}$, by convention, $\Lambda^{0}(\overline{P})=\overline{\mathbbm{k}}$, which is simple. Henceforth we suppose that $t>0$. By construction, $\Lambda^{t}(\overline{P})$ is a direct summand of $\overline{P}^{\otimes t}$. Hence, any composition factor of $\Lambda^{t}(\overline{P})$ is also one of $\overline{P}^{\otimes t}$. Now, Corollary 4.32 implies that any such composition factor is non-zero when evaluated on $\mathbf{t+1}$. Since $\dim\Lambda^{t}(\overline{P})(\mathbf{t+1})=1$ by Lemma 5.6, it follows that there is only one such. ∎
@@
$x_{1}]\otimes\ldots\otimes[x_{t}]\mapsto[x_{1}]\wedge\ldots\wedge[x_{t}].$
@@
For the surjectivity, it suffices to observe that the image of $x_{1}]\otimes([x_{2}]-[x_{1}])\otimes\ldots\otimes([x_{t}]-[x_{1}])$ is $x_{1}]\wedge\ldots\wedge[x_{t}]$. The final statement holds since $\Lambda^{t}(P^{\mathbf{FA}})$ is projective. ∎
@@
To show that this exhibits $\Lambda^{t}(P^{\mathbf{FA}})$ as the projective cover of the simple functor $\Lambda^{t-1}(\overline{P})$, it suffices to show that $\Lambda^{t}(P^{\mathbf{FA}})$ is indecomposable . For $t=0$, using Convention 5.5, this is Lemma 3.4, hence we suppose that $t>0$. It suffices to show that $\mathrm{End}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{t}(P^{\mathbf{FA}}))=\mathbbm{k}$. Using the canonical surjection $P^{\mathbf{FA}}_{\mathbf{t}}\twoheadrightarrow\Lambda^{t}(P^{\mathbf{FA}})$, this follows from the equality $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{t}},\Lambda^{t}(P^{\mathbf{FA}}))=\mathbbm{k}$ given by Yoneda’s lemma and the fact that that $\Lambda^{t}(\mathbbm{k}[\mathbf{t}])\cong\mathbbm{k}$, by Lemma 5.6. ∎
The short exact sequences (5.2) splice to yield the following (a related result occurs in [10]):
@@
$\ldots\rightarrow\Lambda^{t}(P^{\mathbf{FA}})\rightarrow\Lambda^{t-1}(P^{\mathbf{FA}})\rightarrow\ldots\rightarrow\Lambda^{1}(P^{\mathbf{FA}})\rightarrow\Lambda^{0}(P^{\mathbf{FA}})=\mathbbm{k}\rightarrow\mathbbm{k}_{\mathbf{0}}\rightarrow 0.$
@@
For $t$ a positive integer, Wiltshire-Gordon [10] defines $D_{t}$ to be the cokernel of the map $\Lambda^{t+1}(P^{\mathbf{FA}})\rightarrow\Lambda^{t}(P^{\mathbf{FA}})$ appearing in Proposition 5.10. By the above identifications, one has $D_{t}\cong\Lambda^{t-1}(\overline{P})$.
@@ 5.3. Classification of the indecomposable projectives in $\mathcal{F}(\mathbf{FA})$
For $n\in\mathbb{N}$, Theorem 5.3 gives that $S_{\lambda}(\overline{P})$ is an indecomposable projective for $\lambda\neq(1^{n})$ and Theorem 9.3 that $\Lambda^{n}(P^{\mathbf{FA}})$ is an indecomposable projective. One checks using Proposition 5.2 that these are pairwise non-isomorphic, in the sense of the uniqueness in the statement. It remains to show that these exhaust the isomorphism classes of indecomposable projectives.
@@
The maximality of $n$ implies that either $Q$ is isomorphic to $\Lambda^{n+1}(P^{\mathbf{FA}})$ or $Q$ is a quotient of $\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P})$ (these possibilities are mutually exclusive). In the second case, we must have $n>0$, since $\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P})$ is zero for $n=0$; then, using the splitting of $\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P})$ into isotypical components, $Q$ must be isomorphic to $S_{\lambda}(\overline{P})$, for a unique $\lambda\vdash n$ such that $\lambda\neq(1^{n})$. ∎
@@
This theorem provides an abstract classification of the simple functors in $\mathcal{F}(\mathbf{FA})$. Such a classification was given by Wiltshire-Gordon in [10], using a slightly different method. The simple functors are made explicit in Section 6 below.
@@ 6. The simple $\mathbbm{k}\mathbf{FA}$-modules
In this section, the simple objects of $\mathcal{F}(\mathbf{FA})$ (aka. $\mathbbm{k}\mathbf{FA}$-modules) are made explicit. These are related to the structure of $\mathbbm{k}{\mathbf{FI}}(-,-)$, thus relating to results of Feigin and Shoikhet [2].
@@ 6.1. The simple functors
Consider the case $n>0$; Proposition 3.13 exhibits $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$ as a quotient of $P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n-1}$. As in the proof of Theorem 5.13, the latter decomposes as $\Lambda^{n}(P^{\mathbf{FA}})\oplus(\overline{P}^{\otimes n}/\Lambda^{n}(\overline{P}))\oplus(\overline{P}^{\otimes n-1}/\Lambda^{n-1}(\overline{P}))$, in which each summand is projective. Moreover, $\overline{P}^{\otimes n-1}/\Lambda^{n-1}(\overline{P})$ is a direct summand of $P^{\mathbf{FA}}_{\mathbf{n-1}}$. Since $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$ is zero when evaluated on $\mathbf{n-1}$, Yoneda’s lemma implies that there are no non-trivial maps from $\overline{P}^{\otimes n-1}/\Lambda^{n-1}(\overline{P})$ to $\mathbbm{k}{\mathbf{FI}}(\mathbf{n},-)$. Hence the surjection of Proposition 3.13 restricts to a surjection
@@
To conclude, it suffices to check that this restriction is not surjective. This is seen by evaluating on $\mathbf{n}$, as follows. We claim that the map $\overline{P}^{\otimes n}(\mathbf{n})\rightarrow\mathbbm{k}\mathfrak{S}_{n}$ maps to the kernel of the sign representation $\mathsf{sgn}\colon\mathbbm{k}\mathfrak{S}_{n}\rightarrow\mathbbm{k}$.
If $n=1$, the result is immediate, since $\overline{P}(\mathbf{1})=0$; for $n>1$, $\overline{P}(\mathbf{n})$ has dimension n1n-1, with basis $\{[i]-[n]\mid 1\leq i\leq n-1\}$, by Lemma 3.3. Thus $\overline{P}^{\otimes n}(\mathbf{n})$ has basis given by elements of the form $\bigotimes_{j=1}^{n}([i_{j}]-[n])$, where $i_{*}:\mathbf{n}\rightarrow\mathbf{n-1}$. It is straightforward to check that, if $i_{*}$ is not surjective, then the associated basis element maps to zero. Hence suppose that $i_{*}$ is surjective; thus there exist $j_{1}<j_{2}$ such that $i_{j_{1}}=i_{j_{2}}$. To illustrate the argument, consider the case $n=2$, so that $j_{1}=1$ and $j_{2}=2$ and one has the (unique) basis element $([1]-[2])\otimes([1]-[2])$; this maps to $\mathrm{Id}]-[\tau]\in\mathbbm{k}\mathfrak{S}_{2}$, which lies in the kernel of the sign representation. The general case follows by extending this analysis. ∎
@@
We first prove that, for $\lambda\neq(1^{n})$, $C_{\lambda}=\tilde{C}_{\lambda}$ is simple. Suppose otherwise, since $C_{\lambda}(\mathbf{n})$ is a simple $\mathfrak{S}_{n}$-module by Lemma 6.5, then $C_{\lambda}$ would have a non-zero composition factor $G$ such that $G(\mathbf{n})=0$. Since $C_{\lambda}$ is a quotient of $\overline{P}^{\otimes n}$ by Corollary 6.3, $G$ is a composition factor of the latter, hence Corollary 4.32 gives that $G(\mathbf{n+1})\neq 0$.
Now, our understanding of the indecomposable projective generators of $\mathcal{F}(\mathbf{FA})$ (see Theorem 5.13) implies that either $G$ is a quotient of $S_{\mu}(\overline{P})$ for some partition $\mu\vdash n+1$ or $G$ is $\Lambda^{n}(\overline{P})$, hence a quotient of $\Lambda^{n+1}(P^{\mathbf{FA}})$. The first case is excluded since $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(S_{\mu}(\overline{P}),\overline{P}^{\otimes n})=0$, as a consequence of Corollary 4.21, since $S_{\mu}(\overline{P})$ is a direct summand of $\overline{P}^{\otimes n+1}$. The second case is excluded similarly, since $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n+1}(P^{\mathbf{FA}}),S_{\lambda}(\overline{P}))=0$ for $\lambda\neq(1^{n})$, again as a consequence of Corollary 4.21; explicitly, any such morphism would factor across $\Lambda^{n}(\overline{P})$ and $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\Lambda^{n}(\overline{P}),S_{\lambda}(\overline{P}))=0$ by Proposition 5.2.
@@
In [9], Sam, Snowden and Tosteson gave a model for the category of $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules using polynomial representations of the Witt Lie algebra. They observed that, under their equivalence (and using duality), the classification of the simple $\mathbbm{k}\mathbf{FA}$-modules given in [10] corresponds to the classification of certain irreducible modules over the Witt Lie algebras given by Feigin and Shoikhet [2].